OpenStudy (anonymous):

integral(x(36x^2+35)^14)dx

OpenStudy (anonymous):

\[\int\limits(x(36x^2+35)^{14} dx\]

OpenStudy (anonymous):

x>0

OpenStudy (anonymous):

\[\int\limits{f'(x)g(x)}dx=f(x)g(x)-\int\limits{f(x)g'(x)}dx\]Let f'(x)=x and f(x)=1/2 x^2. Use the relation above.